I need help with #7In Problems 5-8, find the derivative, but do not simplify your answer. 5. y = (7x6 – 5x4 + 2x2 – 1)(4x + 3x? – 5x2 + 3x) 6.
Answer #1

15. y = (72% - 5x + 2x²-1) (4x² + 3x7_ sx² + 3x) a = 67x -5x4 +92-1) d (4x +387-5x +330) + [4x2° +387-sx?+33) xd (726_5x* +2(اد 22 ) 4 (ر) + (1- s7 (1+ و )= ( ) (5 - 3 ) + (s-3) (1- عود )= الوار (523) + (و۔ (بداو - ) =